OpenStudy (anonymous):

Angelo took a group of people to the movies and spent $80 on tickets. Adult tickets cost $10, and child tickets cost $6. The equation to represent his cost is 10A + 6C = $80. If there were 5 children in the group, how many adult tickets did he purchase? Show your work and use a complete sentence to explain your solution in the context of the original problem.

OpenStudy (anonymous):

@Hero @ParthKohli @ganeshie8 @jigglypuff314

OpenStudy (igreen):

Spent $80 total. Child tickets cost $6 If 5 children, how many adults? 5 \(\times\) 6 = 30 If he spent $30 on children, he must've spent $50 on adults. And 50 / 10 = 5. So there were 5 adults.

OpenStudy (anonymous):

10a + 30 = 80 i got that part before i posted my question

jigglypuff314 (jigglypuff314):

10a + 30 = 80 subtract 30 from both sides then divide both sides by 10